OpenStudy (mellisa):

Use a table to find two consecutive integers between which the solution lies. 3.5=2.1-0.4y Between what two consecutive integers does the solution​ lie? nothing ​(Use a comma to separate answers as​ needed.)

OpenStudy (mellisa):

help @michaelbp

OpenStudy (michaelbp):

are you looking for y?

OpenStudy (mellisa):

Thats the problem I don't know what i need to get

OpenStudy (michaelbp):

well y=-3.5

OpenStudy (mellisa):

but it says Between what two consecutive integers does the solution​ lie? i guess that what i need?

OpenStudy (michaelbp):

Consecutive integers are integers that follow each other in order. They have a difference of 1 between every two numbers. In a set of consecutive integers, the mean and the median are equal. If n is an integer, then n, n+1, and n+2 would be consecutive integers.

OpenStudy (mellisa):

cool 1 and 2 thanks

OpenStudy (mellisa):

um 1 and 2 is wrong

OpenStudy (mellisa):

@Google

OpenStudy (mellisa):

@Google

OpenStudy (google):

give me a second

OpenStudy (google):

I couldn't figure out how to do this one but i found an answer online that might be good -3 | 1.5-.4y = 1.5 - .4 x(-3) = 2.7 -4 | 1.5 - .4 x (-4) = 3.1 -5 | 1.5 - .4 x (-5) = 3.5 -6 | 1.5 - .4 x (-6) = 3.9 -7 | 1.5 - .4 x (-7) = 4.3

OpenStudy (mellisa):

ok i got it